OpenStudy (anonymous):

What is the answer to this: 2.4^(2/3) / 4.6^(2/3)

OpenStudy (anonymous):

do the ^(2/3) just cancel out?

OpenStudy (unklerhaukus):

\[\frac{2.4^{2/3} }{4.6^{2/3}}=\left(\frac{2.4}{4.6}\right)^{2/3}\]

OpenStudy (anonymous):

hmm thats what i thought but my text gives the answer of just 2.4 / 4.6 = 0.521

OpenStudy (anonymous):

thanks i better go back to my lecturer. thanks mate

OpenStudy (unklerhaukus):

im not sure how they got = 0.521.. you can't just cancel the indicies